In the determination of transference numbers of HCl by moving boundary method a 0.10 M solution of HCl was taken in a cell having uniform area of cross section of 1.6 cm2. At the end of experiment the boundary moved by 11.7 cm and 0.22 g of silver was deposited on the cathode of silver coulometer. Calculate the transference numbers of ions of HCl.


1
Expert's answer
2021-03-26T05:36:52-0400

Transport number of Hions = l x a x c / 1000Q (moving boundary method)

l is length of migration = 11.7 cm

a is area of cross-section =1.6 cm2

c is concentration of solution =0.10 M

Q is charge passed = Equivalent of Ag deposited Q = W/E = 0.22 /108 Faraday

Therefore tH+ = (11.7 x 1.6 x 0.1 x 108) / (0.22 x 1000) = 0.92

Transport number of H+ ions is 0.92

tH+​ + tCl− = 1

tCl− ​= 1− 0.92 = 0.08

Transport number of Cl ions is 0.08.